### v1

# Slices in the loop spaces of symmetric varieties and the formality conjecture

Abstract.
We construct transversal slices to the closures of spherical orbits in loop spaces of symmetric varieties generalizing a construction of Mars-Springer in the finite dimensional setting. We show that those slices are isomorphic to certain Lagrangian subvarieties of the affine Grassmannian slices. We provide applications to the formality conjecture in relative Langlands duality.
1. Introduction
Let $G$ be a complex connected reductive group and let $X=G/K$ be a symmetric variety, where $K$ is the fixed point subgroup of an involution of $G$. Let $B$ be a Borel subgroup of $G$. In [MS], Mars-Springer constructed conical transversal slices to the closure of $B$-orbits on $X$ and used them to show that the $\mathrm{IC}$-complexes for the orbit closures are pointwise pure. This is an important geometric ingredient in their work on Hecke algebra representations associated to symmetric varieties providing a more geometric approach to the results of Lusztig-Vogan in [LV].
In this paper we generalize Mars-Springer’s construction to the setting of loop space LX=X((t))LX=X(\!(t)\!) of $X$. Let G[[t]]G[\![t]\!] be the arc group of $G$. We construct conical transversal slices $S^{\lambda}_{\leq\mu}$ to the closures of G[[t]]G[\![t]\!]-orbits on $LX$ and establish several basic geometric properties of them. We show that the slices $S^{\lambda}_{\leq\mu}$ are isomorphic to the fixed point subvarieties of certain involutions on the affine Grassmannian slices $W^{\lambda}_{\leq\mu}$ for $G$. As an application, we show that $S^{\lambda}_{\leq\mu}$ are Lagrangian subvarieties for $W^{\lambda}_{\leq\mu}$.
One of the motivation for constructing the slices $S^{\lambda}_{\leq\mu}$ is the formality conjecture in relative Langlands duality. In [BZSV], the authors conjecture a version of derived Satake equivalence for symmetric varieties.  [^1] In loc. cit., they consider a more general setting of spherical varieties. Their conjecture implies that the so called de-equivariantized Ext algebra $A_{X}$ for the Satake category of $X$ is formal. A general approach to proving formality is to prove the purity of the Ext algebra $H^{\bullet}(A_{X})$. We carry out this approach when $X$ are classical symmetric varieties. The transversal slice $S^{\lambda}_{\leq\mu}$ is a key ingredient in the proof. As an application, we show that the the de-equivariantized Ext algebra $A_{\mathbb{R}}$ for Satake categories of the real and quaternionic linear groups is formal.
Acknowledgement
The research of T.-H. Chen is supported by NSF grant DMS-2143722.
2. Symmetric varieties
Let $k=\mathbb{C}$ and let $G/k$ a connected reductive group with Lie algebra $\mathfrak{g}=\mathrm{Lie}(G)$. Let $\theta$ be an involution on $G$, and let $K=G^{\theta}$ be the associated symmetric subgroup. The tangent map $d\theta:\mathfrak{g}\rightarrow\mathfrak{g}$ has eigenspace decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$, where $\mathfrak{k}=\mathrm{Lie}(K)$ and $\mathfrak{p}$ is the (1)(-1)-eigenspace. The quotient scheme $X=G/K$ is a symmetric variety. Let $\tau:G\to G,\tau(g)=g\theta(g)^{-1}$. It induces a $G$-equivariant isomorphism $\tau:X\cong G^{\operatorname{inv}\circ\theta,\circ}$ (denoted again by $\tau$) between $X$ and the identity component of the subvariety of $G$ on which $\theta$ acts as inverse. Here $G$ acts on $G^{\operatorname{inv}\circ\theta,\circ}$ by the twisted conjugation action $g\cdot x=gx\theta(g)^{-1}$.
Let $G_{\mathbb{R}}$ be the real form associated to $\theta$, $S_{\mathbb{R}}$ (resp. $T_{\mathbb{R}}$) a $\theta$-stable maximal $\mathbb{R}$-split torus (resp. maximal torus that contains $S_{\mathbb{R}}$) of $G_{\mathbb{R}}$, and $S\subset G$ (resp. $T$) the complexification of $S_{\mathbb{R}}$ and $T_{\mathbb{R}}$. Then $S\subset G^{\operatorname{inv}\circ\theta,\circ}\cong X$. We fix a Borel subgroup $B$ containing $T$.
3. Spherical orbits
[…]
Throughout this paper, we will be concerned exclusively with the topology of loop spaces and related moduli and ignore their potentially non-reduced structure.
We record some results on $G(\mathcal{O})$-orbits on the loop space $LX$, $K(F)$-orbits on $\mathrm{Gr}$ (via the right $LK$-action) in [NadlerMatsuki]. We have maps
$\pi_{*}:\pi_{1}(G)\stackrel{{\scriptstyle}}{{\to}}\pi_{1}(X)\leftarrow\Lambda_{S}^{+}:[-]$
where the right arrow assigns to a loop its homotopy class. Let $\mathcal{L}\subset\Lambda_{S}^{+}$ be the inverse image of [][-] of the image of $\pi_{*}$. Note that $\mathcal{L}=\Lambda_{S}^{+}$ if $K$ is connected.
We order $\mathcal{L}$ by restricting the natural order from the dominant coweights: $\lambda\leq\mu$ if and only if $\mu-\lambda$ is a non-negative sum of positive roots of $G$.
Proposition 1.
[NadlerMatsuki, Theorem 4.2, Theorem 4.3, Theorem 10.1]
  • (i)
    The $G(\mathcal{O})$-orbits on $LX$ are represented by $\Lambda_{S}^{+}$: for each $\mu\in\Lambda_{S}^{+}$, there exists $x_{\mu}\in LX$ such that $\tau(x_{\mu})=t^{\mu}\in LS$ and $LX_{\mu}=G(\mathcal{O})\cdot x_{\mu}$ is the corresponding orbit. Let $\overline{LX}_{\mu}$ be the closure of $LX_{\mu}$. We have $\overline{LX}_{\mu}\subset\bigsqcup_{\lambda\leq\mu}LX_{\lambda}$.
  • (ii)
    The $LK$-orbits on $\mathrm{Gr}$ are parametrized by $\mathcal{L}$: for each $\mu\in\mathcal{L}$, there exists $y_{\mu}\in\mathrm{Gr}$ and a lift $\tilde{y}_{\mu}\in LG$ such that $\tau(\tilde{y}_{\mu})=t^{\mu}\in LS$ and $\mathrm{Gr}^{X}_{\mu}=y_{\mu}\cdot LK$ is the corresponding orbit. Let $\overline{\mathrm{Gr}}^{X}_{\mu}$ be the closure of $\mathrm{Gr}_{\mu}^{X}$. We have $\overline{\mathrm{Gr}}^{X}_{\mu}=\bigsqcup_{\lambda\leq\mu}{\mathrm{Gr}}^{X}_{\lambda}$.
  • (iii)
    The union of spherical orbits $LX_{\mathcal{L}}:=\bigcup_{\lambda\in\mathcal{L}}LX_{\lambda}$ is a union of connected component of $LX$. The image of the natural map $\pi:LG\to LX$ is equal to $LX_{\mathcal{L}}$ and induces an isomorphism $LG/LK\cong LX_{\mathcal{L}}$. Let $p:LG\to\mathrm{Gr}$ be the quotient map. Then for any $\mu\in\mathcal{L}$ we have $p^{-1}(\mathrm{Gr}^{X}_{\mu})=\pi^{-1}(LX_{\mu})$.
  • (iv)
    Let $\mu\in\mathcal{L}$. We have $\overline{LX}_{\mu}=\bigsqcup_{\lambda\leq\mu,\lambda\in\mathcal{L}}LX_{\lambda}$
Remark 2.
It is interesting to note that the proof of the orbits closure relation in (ii) makes use of real affine Grassmannians. Thus these results are not available at the moment over positive characteristic base fields.
4. Transversal slices
In this section, we construct transversal slices of $G(\mathcal{O})$-obits on the loop space $LX$ of a symmetric variety $X$.
4.1. The construction
Our construction of slices of $G(\mathcal{O})$-orbits on $LX$ is a variation of the construction of slices of $B$-orbtis on the symmetric variety $X$ due to Mars and Springer [MS, §6]. Let $\nu\in\Lambda_{S}^{+}$ and let $LX_{\nu}=G(\mathcal{O})\cdot x_{\nu}$ be the corresponding orbit. Since $\theta(t^{\nu})=t^{-\nu}$, the map $\psi_{\nu}=\mathrm{Ad}_{t^{\nu}}\circ\theta:LG\to LG$ is an involution on $LG$ with fixed points $(LG)^{\psi_{\nu}}=\mathrm{Ad}_{x_{\nu}}LK$. Let $L^{-}G=G[t^{-1}]$, $L^{<0}G=\ker(L^{-}G\rightarrow G)$, and $L^{-}\mathfrak{g},L^{<0}\mathfrak{g}$ their Lie algebras. Note that $L^{-}G,L^{<0}G$ are $\theta$-stable. We define
$S^{\nu}:=L^{<0}G\cap\psi_{\nu}L^{<0}G\cdot x_{\nu}\subset LX.$
Note that we have canonical isomorphism
$S^{\nu}\cong L^{<0}G\cap\psi_{\nu}L^{<0}G/(L^{<0}G)^{\psi_{\nu}}.$
The following proposition shows that $S^{\nu}$ is a transversal slice to the orbit $LX_{\nu}$ inside $LX$.
Theorem 3.
  • (i)
    $S^{\nu}\subset LX$ is transversal to $LX_{\nu}=G(\mathcal{O})\cdot x_{\nu}$ at $x_{\nu}$.
  • (ii)
    Let $r_{s}:t\mapsto s^{-2}t$ be rotation automorphism of $LG$, $s\in\mathbb{G}_{m}$. Then $\phi_{s}:(x\mapsto s^{\nu}r_{s}(x))$ defines a $\mathbb{G}_{m}$-action on $S^{\nu}$, which fixes $x_{\nu}$, and contracts $S^{\nu}$ to $x_{\nu}$.
  • (iii)
    The multiplication map $G(\mathcal{O})\times S^{\nu}\rightarrow LX$ is submersion, that is, the induced maps on tangent spaces are surjective.
Proof.
(i): The transversality means we have direct sum of tangent spaces at $x_{\nu}$:
(1) $T_{x_{\nu}}LX=T_{x_{\nu}}S^{\nu}\oplus T_{x_{\nu}}LX_{\nu}.$
Explicitly, this is
(2) $L\mathfrak{g}/(L\mathfrak{g})^{\psi_{\nu}}=L^{<0}\mathfrak{g}\cap\psi_{\nu}(L^{<0}\mathfrak{g})/(L^{<0}\mathfrak{g}\cap\psi_{\nu}(L^{<0}\mathfrak{g}))^{\psi_{\nu}}\oplus L^{+}\mathfrak{g}/(L^{+}\mathfrak{g})^{\psi_{\nu}}.$
First we show the subspaces on RHS span the LHS. We need to show
(3) $L^{<0}\mathfrak{g}\subset L^{<0}\mathfrak{g}\cap\psi_{\nu}(L^{<0}\mathfrak{g})+(L\mathfrak{g})^{\psi_{\nu}}+L^{+}\mathfrak{g}.$
Recall $T$ is $\theta$-stable, thus $LT$ is stable under $\psi_{\nu}$. Let $\mathfrak{t}=\mathrm{Lie}(T)$, then $L^{<0}\mathfrak{t}\subset L^{<0}\mathfrak{g}\cap\psi_{\nu}(L^{<0}\mathfrak{g})$. So $L\mathfrak{t}$ is contained in the RHS of (3).
It remains to show that for any affine root $\tilde{\alpha}=\alpha+m\delta$, where $\alpha$ is a root, $m<0$, $\delta$ is the imaginary root, the affine root subspace $\mathfrak{g}_{\tilde{\alpha}}=t^{m}\mathfrak{g}_{\alpha}$ is contained in the RHS. Recall that $\psi_{\nu}$ preserves $LT$, thus it acts on the set of affine root.
If $\psi_{\nu}(\tilde{\alpha})=\tilde{\alpha}$, then $\mathfrak{g}_{\tilde{\alpha}}\subset L^{<0}\mathfrak{g}\cap\psi_{\nu}(L^{<0}\mathfrak{g})$.
If $\psi_{\nu}(\tilde{\alpha})\neq\tilde{\alpha}$, then $\psi_{\nu}$ acts on the two-dimensional space $\mathfrak{g}_{\tilde{\alpha}}\oplus\mathfrak{g}_{\psi_{\nu}(\tilde{\alpha})}$ which doesn’t preserve the two summands. When $\mathfrak{g}_{\psi_{\nu}(\tilde{\alpha})}\subset L^{<0}\mathfrak{g}$, applying $\psi_{\nu}$ we get $\mathfrak{g}_{\tilde{\alpha}}\subset L^{<0}\mathfrak{g}\cap\psi_{\nu}(L^{<0}\mathfrak{g})$. Otherwise, $\mathfrak{g}_{\psi_{\nu}(\tilde{\alpha})}\subset L^{+}\mathfrak{g}$. We have surjection
$(\mathfrak{g}_{\tilde{\alpha}}\oplus\mathfrak{g}_{\psi_{\nu}(\tilde{\alpha})})^{\psi_{\nu}}\twoheadrightarrow\mathfrak{g}_{\tilde{\alpha}}\mod L^{+}\mathfrak{g}.$
Thus $\mathfrak{g}_{\tilde{\alpha}}\subset(L\mathfrak{g})^{\psi_{\nu}}+L^{+}\mathfrak{g}$.
Next, we show that the RHS of (2) is a direct sum. Let $u\in L^{<0}\mathfrak{g}\cap\psi_{\nu}(L^{<0}\mathfrak{g}),v\in L^{+}\mathfrak{g}$, such that $u+v\in(L\mathfrak{g})^{\psi_{\nu}}$. We want to show $u,v\in(L\mathfrak{g})^{\psi_{\nu}}$. Note that
$L^{+}\mathfrak{g}=L^{+}\mathfrak{g}\cap\psi_{\nu}(L^{<0}\mathfrak{g})\oplus L^{+}\mathfrak{g}\cap\psi_{\nu}(L^{+}\mathfrak{g}).$
Decompose $v=v_{1}+v_{2}$ as above. We get
$u+(v_{1}+v_{2})=\psi_{\nu}(u+v_{1}+v_{2})=(\psi_{\nu}(u)+\psi_{\nu}(v_{1}))+\psi_{\nu}(v_{2}),$
where $\psi_{\nu}(u)+\psi_{\nu}(v_{1})\in L^{<0}\mathfrak{g}$ and $\psi_{\nu}(v_{2})\in L^{+}\mathfrak{g}$. We obtain
$u=\psi_{\nu}(u)+\psi_{\nu}(v_{1}),\qquad v_{1}+v_{2}=\psi_{\nu}(v_{2}).$
Thus
$\psi_{\nu}(v_{1})=u-\psi_{\nu}(u)=\psi_{\nu}(\psi_{\nu}(v_{2})-v_{2})=v_{2}-\psi_{\nu}(v_{2})\in L^{<0}\mathfrak{g}\cap L^{+}\mathfrak{g}=0.$
So $v_{1}=0$, $u=\psi_{\nu}(u)\in(L\mathfrak{g})^{\psi_{\nu}}$, $v=v_{2}=\psi_{\nu}(v_{2})\in(L\mathfrak{g})^{\psi_{\nu}}$.
(ii): We have $\tau(s^{\nu}r_{s}(x_{\nu}))=s^{\nu}\tau(r_{s}(x_{\nu}))\theta(s^{\nu})^{-1}=s^{\nu}r_{s}(\tau(x_{\nu}))s^{\nu}=s^{\nu}(s^{-2}t)^{\nu}s^{\nu}=t^{\nu}$. Thus $\phi_{s}(x_{\nu})=x_{\nu}\in LX$.
For any $gx_{\nu}\in S^{\nu}$ where $g\in L^{<0}G\cap\psi_{\nu}L^{<0}G=L^{<0}G\cap\mathrm{Ad}_{t^{\nu}}L^{<0}G$, since $s^{\nu}r_{s}(x_{\nu})=x_{\nu}$, we have
$\phi_{s}(gx_{\nu})=s^{\nu}r_{s}(g)r_{s}(x_{\nu})=(\mathrm{Ad}_{s^{\nu}}r_{s}(g))x_{\nu}.$
Since $s^{\nu}\in G(k)$ normalizes $L^{<0}G$, we see $\mathrm{Ad}_{s^{\nu}}r_{s}(g)\in L^{<0}G\cap\mathrm{Ad}_{t^{\nu}}L^{<0}G$. Thus $\phi_{s}$ preserves $S^{\nu}$.
Finally, in the above, write $g=\mathrm{Ad}_{t^{\nu}}h$, $h\in L^{<0}G$. Then
$(\mathrm{Ad}_{s^{\nu}}r_{s}(g))x_{\nu}=(\mathrm{Ad}_{(ts^{-1})^{\nu}}r_{s}(h))x_{\nu}=r_{s^{1/2}}(\mathrm{Ad}_{t^{\nu}}r_{s^{1/2}}(h))x_{\nu},$
which goes to $x_{\nu}$ as $s$ goes to 0.
(iii): It suffices to show that the multiplication
$G(\mathcal{O})\times(L^{<0}G\cap\mathrm{Ad}_{t^{\nu}}L^{<0}G)x_{\nu}\rightarrow LG/(LG)^{\psi_{\nu}}$
is submersion. At an arbitrary point $(g,ux_{\nu})$, the tangent map is
$L^{+}\mathfrak{g}\times(L^{<0}\mathfrak{g}\cap\mathrm{Ad}_{t^{\nu}}L^{<0}\mathfrak{g}/\mathrm{Ad}_{u}(L^{<0}\mathfrak{g})^{\psi_{\nu}}\cap L^{<0}\mathfrak{g}\cap\mathrm{Ad}_{t^{\nu}}L^{<0}\mathfrak{g})$ $\rightarrow L\mathfrak{g}/\mathrm{Ad}_{gu}(L\mathfrak{g})^{\psi_{\nu}},$
$(x,y)$ $\mapsto x+\mathrm{Ad}_{g}y.$
We need to show:
$L^{+}\mathfrak{g}+\mathrm{Ad}_{g}(L^{<0}\mathfrak{g}\cap\mathrm{Ad}_{t^{\nu}}L^{<0}\mathfrak{g})+\mathrm{Ad}_{gu}(L\mathfrak{g})^{\psi_{\nu}}=L\mathfrak{g}.$
Conjugate by $(gu)^{-1}$, this becomes
(4) $\mathrm{Ad}_{u^{-1}}L^{+}\mathfrak{g}+L^{<0}\mathfrak{g}\cap\mathrm{Ad}_{t^{\nu}}L^{<0}\mathfrak{g}+(L\mathfrak{g})^{\psi_{\nu}}=L\mathfrak{g}.$
Note that when $u=1$, (4) becomes (3). For general $u$, note that the $\mathbb{G}_{m}$-action $\phi_{s}$ in (ii) preserve each summand in the above except $u$, where $\lim_{s\rightarrow 0}\phi_{s}(u)=1$. Thus it suffices to prove (4) with $u$ replaced by $\phi_{s}(u)$ where $s$ is close to 0. Assume $\phi_{s}(u),\phi_{s}(u^{-1})\in G(t^{-1}k+t^{-2}k+\cdots+t^{-r-1}k)\subset L^{<0}G$, where $r\in\mathbb{N}$ is independent of $s$. Then $\mathrm{Ad}_{\phi_{s}(u)}t^{2r}L^{+}\mathfrak{g}\subset L^{+}\mathfrak{g}$, $t^{2r}L^{+}\mathfrak{g}\subset\mathrm{Ad}_{\phi_{s}(u^{-1})}L^{+}\mathfrak{g}$. Combining this with (3), it suffices to show that a finite set of vectors in $\mathrm{Ad}_{\phi_{s}(u^{-1})}L^{+}\mathfrak{g}/t^{2r}L^{+}\mathfrak{g}$ are linearly independent. Since the coefficients of these vectors with respect to a basis of affine root subspaces are polynomials in $s$, we can deduce the linear independence for small $s$ from the case of $s=0$, which is known from (3). This completes the proof. ∎
Remark 4.
Note that the definition of the slice $S^{\nu}$ makes senses for general spherical varieties $X$ but it might not be a transversal slices to $G(\mathcal{O})$-orbits on $LX$. The proof above uses the involution $\theta$ (or rather $\psi_{\lambda}$) in an essential way.
4.2. Slice intersection with orbits
For any $\lambda,\mu\in\Lambda_{S}^{+}$, let $S^{\lambda}_{\mu}=S^{\lambda}\cap LX_{\mu}$ and $S^{\lambda}_{\leq\mu}=S^{\lambda}\cap\overline{LX}_{\mu}$.
Corollary 5.
If $S^{\lambda}_{\mu}\neq\varnothing$ then $\lambda\leq\mu$. Moreover, when $\lambda,\mu\in\mathcal{L}$, $S^{\lambda}_{\mu}\neq\varnothing$ if and only if $\lambda\leq\mu$.
Proof.
Assume $S^{\lambda}_{\mu}=S^{\lambda}\cap LX_{\mu}\neq\varnothing$. Recall that we have $\mathbb{G}_{m}$-action $\phi_{s}$(depended on $\lambda$) on $S^{\lambda}$. This action also acts on all the $G(\mathcal{O})$-orbits $LX_{\mu}$. The contraction of $S^{\lambda}\cap LX_{\mu}$, which gives $x_{\lambda}$, must be contained in $\overline{LX}_{\mu}$. Thus $\lambda\leq\mu$.
Assume $\lambda\leq\mu\in\mathcal{L}$, then by Proposition 1.(iii) we have $x_{\lambda}\in S^{\lambda}\cap\overline{LX}_{\mu}$. Proposition 3.(iii) implies that the morphism $G(\mathcal{O})\times(S^{\lambda}\cap\overline{LX}_{\mu})\rightarrow\overline{LX}_{\mu}$ has open image. Therefore the image has nonempty intersection with $LX_{\mu}$. ∎
By Proposition 3 and Corollary 5, we immediately obtain
Corollary 6.
Let $\lambda\leq\mu\in\mathcal{L}$.
  • (i)
    $S^{\lambda}_{\leq\mu}\subset\overline{LX}_{\mu}$ is transversal to $LX_{\lambda}$ at $x_{\lambda}$.
  • (ii)
    Let $r_{s}:t\mapsto s^{-2}t$ be rotation automorphism of $LG$, $s\in\mathbb{G}_{m}$. Then $\phi_{s}:(x\mapsto s^{\lambda}r_{s}(x))$ defines a $\mathbb{G}_{m}$-action on $S^{\lambda}_{\leq\mu}$, which fixes $x_{\lambda}$, and contracts $S^{\lambda}$ to $x_{\lambda}$.
  • (iii)
    The multiplication map $G(\mathcal{O})\times S^{\lambda}_{\leq\mu}\rightarrow\overline{LX}_{\mu}$ is a submersion.
4.3. Relation to affine Grassmannian slices
For any $\lambda\in\Lambda_{T}^{+}$, we write $LG_{\lambda}=L^{+}Gt^{\lambda}L^{+}G$ for the $L^{+}G\times L^{+}G$-orbit of $t^{\lambda}$ in $LG$, and $\overline{LG_{\lambda}}$ its closure. The affine Grassmannian slice of $LG_{\lambda}$ in $LG$ is defined as
$W^{\lambda}=(L^{<0}G\cap\mathrm{Ad}_{t^{\lambda}}L^{<0}G)t^{\lambda}.$
For any pair $\lambda,\mu\in\Lambda_{T}^{+}$, we define
$W^{\lambda}_{\mu}=W^{\lambda}\cap LG_{\mu},\ \ \ \ \ W^{\lambda}_{\leq\mu}=W^{\lambda}\cap\overline{LG_{\mu}}.$
We have the following well-known facts, see for example [ZhuIntroGr]:
Proposition 7.
  • (i)
    $W^{\lambda}_{\mu}$ is non-empty if and only if $\lambda\leq\mu$.
  • (ii)
    $W^{\lambda}_{\mu}$ is smooth and connected (if non-empty)
  • (iii)
    $W^{\lambda}_{\mu}$ is finite type affine and normal
  • (iv)
    Assume $k=\mathbb{C}$, we have $\dim W^{\lambda}_{\mu}=W^{\lambda}_{\leq\mu}=\langle 2\rho,\mu-\lambda\rangle$.
Consider the map $\operatorname{inv}\circ\theta:LG\to LG$ sending $g\to\theta(g)^{-1}$ and the map $\rho_{\lambda}(s):LG\to LG$ sending $\gamma(t)\to\rho_{\lambda}(s)(\gamma(t))=s^{\lambda}\gamma(s^{-2}t)s^{\lambda}$.
Lemma 8.
Let $\lambda\leq\mu\in\Lambda_{S}^{+}$.
  • (i)
    the two maps $\rho_{\lambda}(s)$ and $\operatorname{inv}\circ\theta$ commute with each other.
  • (ii)
    The map $\operatorname{inv}\circ\theta:LG\to LG$ restricts to an involution on $W^{\lambda}$, $W^{\lambda}_{\mu}$, and $W^{\lambda}_{\leq\mu}$.
  • (iii)
    The map $\rho_{\lambda}(s)$ restricts to a $\mathbb{G}_{m}$-action $\rho_{s}$ on $W^{\lambda},W^{\lambda}_{\mu},W^{\lambda}_{\leq\mu}$. The $\mathbb{G}_{m}$-action $\rho_{s}$ on $W^{\lambda}$ and $W^{\lambda}_{\leq\mu}$ is contrating with unique fixed point $t^{\lambda}$.
Proof.
Since the proof is a direct computation, we only give a proof of (iii). We have $\operatorname{inv}\circ\theta(t^{\lambda})=t^{\lambda}$, and hence for any $\gamma(t)t^{\lambda}\in W^{\lambda}$, we have
$\rho_{\lambda}(s)(\gamma(t)t^{\lambda})=s^{\lambda}\gamma(s^{-2}t)(s^{-2}t)^{\lambda}s^{\lambda}=s^{\lambda}\gamma(s^{-2}t)s^{-\lambda}t^{\lambda}\in W^{\lambda}$
Thus $\rho_{\lambda}(s)$ preserves $W^{\lambda}$. Since $\gamma\in L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G$, we have $s^{\lambda}\gamma(s^{-2}t)s^{-\lambda}$ goes to the unit $e\in LG$ as $s$ goes to 0, and it follows that the $\mathbb{G}_{m}$-action on $W^{\lambda}$ is contracting with unique fixed point $t^{\lambda}$. Since $LG_{\mu}$ is obviously stable under $\rho_{\lambda}(s)$, the $\mathbb{G}_{m}$-action preserves $W^{\lambda}_{\mu}$ and $W^{\lambda}_{\leq\mu}$. The desired claim follows. ∎
Consider the fixed points subvarieties
(5) $L^{\lambda}=(W^{\lambda})^{\operatorname{inv}\circ\theta},\ \ \ L^{\lambda}_{\mu}=(W_{\mu}^{\lambda})^{\operatorname{inv}\circ\theta},\ \ \ \ L^{\lambda}_{\leq\mu}=(W^{\lambda}_{\leq\mu})^{\operatorname{inv}\circ\theta}.$
Part (i) of above lemma implies that those fixed points subvarieties are stable under the $\mathbb{G}_{m}$-action $\rho_{s}$. Moreover since $t^{\lambda}\in L^{\lambda}$, the $\mathbb{G}_{m}$-action on $L^{\lambda}$ and $L^{\lambda}_{\leq\mu}$ is contracting with unique fixed point $t^{\lambda}$. In particular, we see that $L^{\lambda}_{\mu}$ is smooth and $L^{\lambda}_{\leq\mu}$ is affine and connected.
Recall the embedding
(6) $\tau:LX\rightarrow(LG)^{\operatorname{inv}\circ\theta}\subset LG,\ \ \ \ \tau(\gamma)=\gamma\theta(\gamma)^{-1}.$
Proposition 9.
For any $\lambda,\mu\in\Lambda_{S}^{+}$,
  • (i)
    the map $\tau:LX\to LG$ induces $\mathbb{G}_{m}$-equivariant isomorphisms
    $S^{\lambda}\simeq L^{\lambda},\ \ \ \ S^{\lambda}_{\mu}\simeq L^{\lambda}_{\mu},\ \ \ \ S^{\lambda}_{\leq\mu}\simeq L^{\lambda}_{\leq\mu}.$
  • (ii)
    $S^{\lambda}_{\mu}$ (resp. $L^{\lambda}_{\mu}$) is smooth and $S^{\lambda}_{\leq\mu}$ (resp. $L^{\lambda}_{\leq\mu}$) is finite type affine and connected.
Proof.
(i): For any $s\in\mathbb{G}_{m}$ and $\gamma=\gamma(t)\in LX$, we have
$\tau(\phi_{s}(\gamma))=\tau(s^{\lambda}\gamma(s^{-2}t))=s^{\lambda}\tau(\gamma(s^{-2}t))s^{\lambda}=\rho_{\lambda}(s)(\tau(\gamma)),$
and hence $\tau$ is $\mathbb{G}_{m}$-equivariant.
Let $\gamma x_{\lambda}\in S^{\lambda}=L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G\cdot x_{\lambda}$ with $\gamma\in L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G$. Note that
$\tau(x_{\lambda})=t^{\lambda},\ \ \ \ \ t^{\lambda}\theta(\gamma)^{-1}t^{-\lambda}\in L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G,$
and it follow that
$\tau(\gamma x_{\lambda})=\gamma t^{\lambda}\theta(\gamma)^{-1}=(\gamma t^{\lambda}\theta(\gamma)^{-1}t^{-\lambda})t^{\lambda}\in(L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G)t^{\lambda}=W^{\lambda}.$
Since $\tau(LX)\subset(LG)^{\operatorname{inv}\circ\theta}$ and $\tau(LX_{\mu})\subset LG_{\mu}$, we conclude that the map $\tau$ induces embeddings
$\tau^{\lambda}=\tau|_{S^{\lambda}}:S^{\lambda}\to W^{\lambda}\cap(LG)^{\operatorname{inv}\circ\theta}=L^{\lambda},$
$\tau^{\lambda}_{\mu}=\tau^{\lambda}|_{S^{\lambda}_{\mu}}:S^{\lambda}_{\mu}=S^{\lambda}\cap LX_{\mu}\longrightarrow L^{\lambda}\cap LG_{\mu}=L^{\lambda}_{\mu},$
$\tau^{\lambda}_{\leq\mu}=\tau^{\lambda}|_{S^{\lambda}_{\mu}}:S^{\lambda}_{\mu}=S^{\lambda}\cap\overline{LX_{\mu}}\longrightarrow L^{\lambda}\cap\overline{LG_{\mu}}=L^{\lambda}_{\leq\mu}.$
We need to show that $\tau^{\lambda}$, $\tau^{\lambda}_{\mu}$, $\tau^{\lambda}_{\leq\mu}$ are surjective. Since $S^{\lambda}=\cup_{\lambda\leq\mu}S^{\lambda}_{\mu}$, $L^{\lambda}=\cup_{\lambda\leq\mu}L^{\lambda}_{\mu}$, it suffices to show the map $\tau^{\lambda}_{\leq\mu}$ is surjective. We claim that $\tau^{\lambda}_{\leq\mu}$ is smooth and hence the image contains an open subset of the fixed point $t^{\lambda}\in L^{\lambda}_{\leq\mu}$. We conclude that $\tau^{\lambda}_{\leq\mu}$ is surjective by noting that it is $\mathbb{G}_{m}$-equivariant and the $\mathbb{G}_{m}$-action on $L^{\lambda}_{\leq\mu}$ is contrating. To prove the claim, we observe that $S^{\lambda}_{\leq\mu}$ and $L^{\lambda}_{\leq\mu}$ are finite type and hence it suffices to show that $\tau^{\lambda}_{\leq\mu}$ is formally smooth. Since $\tau^{\lambda}_{\leq\mu}$ is the base change of $\tau^{\lambda}:S^{\lambda}\to L^{\lambda}$, it suffices to show $\tau^{\lambda}$ is formally smooth. For this it suffices to show that tangent map
$d\tau^{\lambda}|_{x_{\lambda}}:T_{x_{\lambda}}S^{\lambda}\to T_{t^{\lambda}}L^{\lambda}$
is surjective. We have following identification
${T_{x_{\lambda}}S^{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }$$\scriptstyle{d\tau^{\lambda}|_{x_{\lambda}}}$${(L^{<0}\mathfrak{g}\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}\mathfrak{g})/(L^{<0}\mathfrak{g})^{\psi_{\lambda}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${T_{t^{\lambda}}L^{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }$${(L^{<0}\mathfrak{g}\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}\mathfrak{g})^{\psi_{\lambda}=-1}}$
[…]
Part (ii) follows from (i) and the fact that the assertions are true for $L^{\lambda}_{\mu}$ and $L^{\lambda}_{\leq\mu}$. ∎
Remark 10.
Unlike the case of $W^{\lambda}_{\mu}$ and $W^{\lambda}_{\leq\mu}$, the fixed points $L^{\lambda}_{\mu}\simeq S^{\lambda}_{\mu}$ might be disconnected and $L^{\lambda}_{\leq\mu}\simeq S^{\lambda}_{\leq\mu}$ might be reducible and non-normal (see Remark 18).
4.4. Slices of $LK$-orbits on $\mathrm{Gr}$
For any $\lambda\in\mathcal{L}$ let $\mathrm{Gr}_{\lambda}^{X}=y_{\lambda}\cdot LK\subset\mathrm{Gr}$ be the corresponding orbit and let $\tilde{y}_{\lambda}\in LG$ be a lift as in Proposition 1. Recall the involution $\psi_{\lambda}=\operatorname{Ad}_{t^{\lambda}}\circ\theta$ on $LG$. Let $(L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G)^{\operatorname{inv}\circ\psi_{\lambda}}$ be the subset consisting of $\gamma\in L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G$ such that $\psi_{\lambda}(\gamma)=\gamma^{-1}$. Let
$\tilde{Y}^{\lambda}:=(L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G)^{\operatorname{inv}\circ\psi_{\lambda}}\cdot\tilde{y}_{\lambda}\subset LG.$
The projection map $p:LG\to\mathrm{Gr}$ restricts to an embedding $p:\tilde{Y}^{\lambda}\hookrightarrow\mathrm{Gr}$ and we define
$Y^{\lambda}:=p(\tilde{Y}^{\lambda})\subset\mathrm{Gr}$
to be the image.
Lemma 11.
There exists a unique homomorphism $\kappa_{\lambda}:\mathbb{G}_{m}\to LK$ such that
$s^{\lambda}\tilde{y}_{\lambda}(s^{-2}t)\kappa_{\lambda}(s)=\tilde{y}_{\lambda}(t).$
Proof.
Note that $\tau(\tilde{y}_{\lambda})=\tau(s^{\lambda}\tilde{y}_{\lambda}(s^{-2}t))=t^{\lambda}$ and hence $s^{\lambda}\tilde{y}_{\lambda}(s^{-2}t)=\tilde{y}_{\lambda}(t)k_{s}$ for a unique $k_{s}\in LK$. For any $s_{1},s_{2}\in\mathbb{G}_{m}$, we have
$s_{1}^{\lambda}s_{2}^{\lambda}\tilde{y}_{\lambda}(s_{1}^{-2}s_{2}^{-2}t)=s_{2}^{\lambda}\tilde{y}_{\lambda}(s_{2}^{-2}t)k_{s_{1}}=\tilde{y}_{\lambda}(t)k_{s_{2}}k_{s_{1}}$
and it follows that $k_{s_{1}s_{2}}=k_{s_{2}}k_{s_{1}}$ and the desired homomorphism is given by $\kappa_{\lambda}(s)=k_{s}^{-1}$
We shall show that $Y^{\lambda}$ defines a transversal slices of $\mathrm{Gr}_{\lambda}^{X}$ at $y_{\lambda}$.
Theorem 12.
Let $\lambda\in\mathcal{L}$.
  • (i)
    There is an isomorphism $Y^{\lambda}\cong S^{\lambda}$.
  • (ii)
    $Y^{\lambda}\subset\mathrm{Gr}$ is transversal to $\mathrm{Gr}^{X}_{\lambda}$ at $y_{\lambda}$.
  • (iii)
    Let $r_{s}:t\mapsto s^{-2}t$ be rotation automorphism of $LG$, $s\in\mathbb{G}_{m}$. Then $\phi_{s}:(x\mapsto r_{s}(x)\kappa_{\lambda}(s))$ defines a $\mathbb{G}_{m}$-action on $Y^{\lambda}$, which fixes $y_{\lambda}$, and contracts $Y^{\lambda}$ to $y_{\lambda}$.
  • (iv)
    The multiplication map $Y^{\lambda}\times LK\rightarrow\mathrm{Gr}$ is a submersion.
Proof.
(i): We claim that the quotient map $\pi:LG\to LX_{\mathcal{L}}$ maps $\tilde{Y}^{\lambda}$ isomorphically onto $S^{\lambda}$. Note that the composed map
$sq:\tilde{Y}^{\lambda}\cong(L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G)^{\operatorname{inv}\circ\psi_{\lambda}}\cdot\tilde{y}_{\lambda}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}S^{\lambda}\stackrel{{\scriptstyle\tau}}{{\cong}}L^{\lambda}\cong(L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G)^{\operatorname{inv}\circ\psi_{\lambda}}\cdot t^{\lambda}$
is equal to the ”square map” sending $\gamma\cdot\tilde{y}_{\lambda}\to\gamma^{2}\cdot t^{\lambda}$. Note that $sq$ is injective and equivaraint with respect to the contracting $\mathbb{G}_{m}$-action $\gamma(t)\to\gamma(s^{-1}t)$ on $(L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G)^{\operatorname{inv}\circ\psi_{\lambda}}$. Since the differential of $sq$ at $\tilde{y}_{\lambda}$ is an isomorphism, the image contains an open neighborhood of $t^{\lambda}$ and it follows that $sq$ is onto. Part (i) follows.
Part (i) implies that there is a Cartesian diagram
${G(\mathcal{O})\times\tilde{Y}^{\lambda}\times LK\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${LG\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G(\mathcal{O})\times S^{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${LX_{\mathcal{L}}\cong LG/LK}$
where the horizontal arrows are the multiplication maps and the vertical maps are the natural projection maps. Now part (ii) and (iv) follows from Theorem 3.
[…]
$\phi_{s}(y)=[\gamma(s^{-2}t)\cdot\tilde{y}_{\lambda}(s^{-2}t)\cdot\kappa_{\lambda}(s)]=[\gamma(s^{-2}t)\cdot s^{-\lambda}\tilde{y}_{\lambda}(t)]=[s^{-\lambda}\cdot\gamma(s^{-2}t)\cdot s^{-\lambda}\tilde{y}_{\lambda}(t)]=[s^{\lambda}\cdot\gamma(s^{-2}t)\cdot s^{-\lambda}\tilde{y}_{\lambda}(t)]$
[…]
$s^{-\lambda}\cdot\gamma(s^{-2}t)\cdot s^{-\lambda}\in(L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G)^{\operatorname{inv}\circ\psi_{\lambda}}$
and
$\operatorname{lim}_{s\to 0}s^{\lambda}\cdot\gamma(s^{-2}t)\cdot s^{-\lambda}\tilde{y}_{\lambda}(t)=\tilde{y}_{\lambda}(t),$
it follows that
$\phi_{s}(y)=[s^{-\lambda}\cdot\gamma(s^{-2}t)\cdot s^{-\lambda}\tilde{y}_{\lambda}(t)]\in[(L^{<0}G\cap\operatorname{Ad}_{t^{\lambda}}L^{<0}G)^{\operatorname{inv}\circ\psi_{\lambda}}\cdot\tilde{y}_{\lambda}(t)]=Y^{\lambda}$
and
$\operatorname{lim}_{s\to 0}\phi_{s}(y)=\operatorname{lim}_{s\to 0}[s^{\lambda}\cdot\gamma(s^{-2}t)\cdot s^{-\lambda}\tilde{y}_{\lambda}(t)]=[\tilde{y}_{\lambda}(t)]=y.$
Part (iii) follows. ∎
Let $Y^{\lambda}_{\mu}=Y^{\lambda}\cap\mathrm{Gr}_{\mu}^{X}$ and $Y^{\lambda}_{\leq\mu}=Y^{\lambda}\cap\overline{\mathrm{Gr}}_{\mu}^{X}$. The following follows from Theorem 12 and Corollary 6.
Corollary 13.
Let $\lambda\leq\mu\in\mathcal{L}$.
  • (i)
    $Y_{\leq\mu}^{\lambda}\subset\overline{\mathrm{Gr}}^{X}_{\mu}$ is transversal to $\mathrm{Gr}^{X}_{\lambda}$ at $y_{\lambda}$.
  • (ii)
    Let $r_{s}:t\mapsto s^{-2}t$ be rotation automorphism of $LG$, $s\in\mathbb{G}_{m}$. Then $\phi_{s}:(x\mapsto r_{s}(x)\kappa_{\lambda}(s))$ defines a $\mathbb{G}_{m}$-action on $Y^{\lambda}_{\leq u}$, which fixes $y_{\lambda}$, and contracts $Y^{\lambda}$ to $y_{\lambda}$.
  • (iii)
    The multiplication map $Y^{\lambda}_{\leq u}\times LK\rightarrow\overline{\mathrm{Gr}}^{X}_{\mu}$ is a submersion.
Corollary 14.
We have $\operatorname{dim}S_{\mu}^{\lambda}=\dim S_{\leq\mu}^{\lambda}=\operatorname{dim}L_{\mu}^{\lambda}=\operatorname{dim}L^{\lambda}_{\leq\mu}=Y^{\lambda}_{\mu}=Y^{\lambda}_{\leq\mu}=\langle\rho,\mu-\lambda\rangle$
Proof.
It suffices to consider the case $Y^{\lambda}_{\leq\mu}$. Let $K_{\mathbb{R}}=G_{\mathbb{R}}\cap K$ be the maximal compact subgroup of $K$ and let $\Omega K_{\mathbb{R}}\subset LK$ be the based loop group of $K_{\mathbb{R}}$. $\Omega K_{\mathbb{R}}$ acts freely on $\mathrm{Gr}$ and it follows from Corollary 13 (iii) that the image of $Y^{\lambda}_{\leq\mu}$ in the quotient $\mathrm{Gr}/\Omega K_{\mathbb{R}}$ defines a transversal slices of $\mathrm{Gr}_{\lambda}^{X}/\Omega K_{\mathbb{R}}$ inside $\overline{\mathrm{Gr}}_{\mu}^{X}/\Omega K_{\mathbb{R}}$. According to [CN2], There are homeomorphisms $\mathrm{Gr}_{\lambda}^{X}/\Omega K_{\mathbb{R}}\cong\mathrm{Gr}_{\mathbb{R},\lambda}$, $\lambda\in\mathcal{L}$, where $\mathrm{Gr}_{\mathbb{R},\lambda}=G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})\cdot[t^{\lambda}]\subset\mathrm{Gr}_{\mathbb{R}}$ is the real spherical orbits through $t^{\lambda}]$ inside the real affine Grassmannian $\mathrm{Gr}_{\mathbb{R}}=G_{\mathbb{R}}(F_{\mathbb{R}})/G(\mathcal{O}_{\mathbb{R}})$ (see Section 5.5), and $\dim_{\mathbb{R}}(\mathrm{Gr}_{\mathbb{R},\lambda})=2\langle\rho,\lambda\rangle$. Thus
$\dim Y^{\lambda}_{\leq\mu}=\mathrm{codim}_{\mathbb{C}}(\mathrm{Gr}_{\mu}^{X}/\Omega K_{\mathbb{R}},\mathrm{Gr}_{\lambda}^{X}/\Omega K_{\mathbb{R}})=\frac{1}{2}\mathrm{codim}_{\mathbb{R}}(\mathrm{Gr}_{G_{\mathbb{R}},\mu},\mathrm{Gr}_{G_{\mathbb{R}},\lambda})=\langle\rho,\mu-\lambda\rangle.$
4.5. Lagrangian slices
Recall the affine Grassmannian slice $W^{\nu}_{\leq\mu}$ has a natural Poisson structure and $W^{\lambda}_{\nu}$ for $\nu\leq\lambda\leq\mu$ are symplectic leaves of dimension $\dim W^{\lambda}_{\nu}=2\langle\rho,\mu-\lambda\rangle$ (see, e.g., [KWWY]). In this subsection, we show $L^{\lambda}_{\mu}\subset W^{\lambda}_{\mu}$ is a Lagrangian subvariety. We will assume $K$ is connected, so that $\tau(LX)=(LG)^{\operatorname{inv}\circ\theta,\circ}$. Only in this subsection, we consider $\mathrm{Gr}=LG/L^{+}G$ for convenience of notations, contrary to the convention $\LG\mathrm{Gr}=L^{+}G\backslash$ LG in other sections.
4.5.1. Poisson structures on the loop group and affine Grassmannian
It is well-known that $(L\mathfrak{g},L^{<0}\mathfrak{g},L^{+}\mathfrak{g})$ forms a Manin triple. Precisely, fix a non-degenerate symmetric bilinear invariant form (,)(\ ,\ ) on $\mathfrak{g}$ that is also $\theta$-invariant. Extend it $F$-linearly to $L\mathfrak{g}=\mathfrak{g}(F)$. Composing with residue map, we get a $\mathbb{C}$-valued inner product on $L\mathfrak{g}$, which we still denote by (,)(\ ,\ ). Then $L^{<0}\mathfrak{g},L^{+}\mathfrak{g}$ are both Lagrange subspaces of $L\mathfrak{g}$ with respect to (,)(\ ,\ ). Fix a basis $e_{i}^{*}$ of $L^{<0}\mathfrak{g}$ and a basis $e_{i}$ of $L^{+}\mathfrak{g}$ such that $(e_{i}^{*},e_{j})=\delta_{ij}$. Define $r$-matrix
$r=\sum_{i}e_{i}^{*}\wedge e_{i}\in\wedge^{2}L\mathfrak{g},$
where the sum is taken in the appropriate sense, see [ZhouSlice, Remark 4.5].
The definition of $r$ is independent of the choice of orthonormal basis. Since $\theta$ acts continuously on $L^{<0}\mathfrak{g}$ and $L^{+}\mathfrak{g}$, we have
$\theta(r)=r.$
Define a bi-vector field $\pi\in\Gamma(LG,\wedge^{2}TLG)$ by
$\pi(g):=R_{g*}r-L_{g*}r,$
where $g\in LG$, $L_{g},R_{g}$ are left and right translations by $g$. Then $\pi$ gives a Poisson bi-vector and defines a Poisson structure on $LG$. Moreover, $L^{<0}G,L^{+}G$ are Poisson subgroups of $LG$.
There are two ways to define a bi-vector field on $\mathrm{Gr}$ from $r$. Let $p:LG\rightarrow\mathrm{Gr}$ be the projection map. Define $\bar{\pi}_{1}\in\Gamma(\mathrm{Gr},\wedge^{2}T\mathrm{Gr})$ by $\bar{\pi}_{1}(g)=d_{g}p(\pi_{g})$.
On the other hand, the action of $LG$ on $\mathrm{Gr}$ defines a map $\kappa:L\mathfrak{g}\rightarrow\Gamma(\mathrm{Gr},T\mathrm{Gr})$. Let $\bar{\pi}_{2}=\kappa(r)\in\Gamma(\mathrm{Gr},\wedge^{2}T\mathrm{Gr})$.
Lemma 15.
We have $\bar{\pi}_{1}=\bar{\pi}_{2}$.
Proof.
At any $g\in LG$, let $\bar{g}=p(g)$. Since $e_{i}\in L^{+}\mathfrak{g}$, we have $p(ge^{te_{i}})=\bar{g}$. Thus $d_{g}p(L_{g*}e_{i})=0$, $d_{g}p(L_{g*}r)=0$, $\bar{\pi}_{1}(g)=dp(R_{g*}r)$. Also, for $a\in L\mathfrak{g}$, $\kappa(a)(\bar{g})$ is represented by curve $p(e^{ta}g)$, i.e. $\kappa(a)(\bar{g})=d_{g}p(R_{g*}a)$. Thus $\bar{\pi}_{1}(\bar{g})=\bar{\pi}_{2}(\bar{g})=d_{g}p(R_{g*}r)$. ∎
Denote $\bar{\pi}=\bar{\pi}_{1}=\bar{\pi}_{2}$. From [LYPoisson, Theorem 2.3], we know $\bar{\pi}$ defines a Poisson structure on $\mathrm{Gr}$, and both $L^{+}G$-orbits $p(LG_{\lambda})=L^{+}G\cdot[t^{\lambda}]$ and $L^{<0}G$-orbits $p(W^{\lambda})=L^{<0}G\cdot[t^{\lambda}]$ are Poisson subvarieties. The intersections of $L^{+}G$-orbits and $L^{<0}G$-orbits, when nonempty, are symplectic leaves, c.f. [LYPoisson, Corollary 2.9], [KWWY, Theorem 2.5].
4.5.2. Lagrangian slices
The projection $p:LG\rightarrow\mathrm{Gr}$ restricts to an embedding on $W^{\lambda}$, thus $p:W^{\lambda}_{\mu}\xrightarrow{\sim}L^{<0}G\cdot[t^{\lambda}]\cap L^{+}G\cdot[t^{\mu}]$. The pullback induces a symplectic structure on $W^{\lambda}_{\mu}$.
Theorem 16.
$L^{\lambda}_{\mu}\subset W^{\lambda}_{\mu}$ is a Lagrangian subvariety.
Proof.
By Proposition 7 and Corollary 14, it suffices to show $L^{\lambda}_{\mu}$ is a coisotropic subspace. This is equivalent to that the ideal of $L^{\lambda}_{\mu}$ in $W^{\lambda}_{\mu}$ is closed under Poisson bracket. For this, it suffices to prove for the embedding $L^{\lambda}\subset W^{\lambda}$, where the Poisson structure on $W^{\lambda}$ is pullback from $p:W^{\lambda}\xrightarrow{\sim}L^{<0}G\cdot[t^{\lambda}]\subset\mathrm{Gr}$.
Denote $\sigma=\operatorname{inv}\circ\theta$, $L^{\lambda}=(W^{\lambda})^{\sigma}$. For any $g\in L^{\lambda}$, $g\in(LG)^{\sigma}=\tau(LX)$, so we can write $g=h\theta(h)^{-1}$ for some $h\in LG$. For $x\in L\mathfrak{g}$, if $e^{tx}g\in(LG)^{\sigma}$, then $x=-\mathrm{Ad}_{g}(d\theta(x))$, $d\theta(\mathrm{Ad}_{h^{-1}}x)=-\mathrm{Ad}_{h^{-1}}x$, $x\in\mathrm{Ad}_{h}L\mathfrak{p}$. Thus
$T_{g}(L^{\lambda})=R_{g*}(L^{<0}\mathfrak{g}\cap\mathrm{Ad}_{t^{\lambda}}L^{<0}\mathfrak{g}\cap\mathrm{Ad}_{h}L\mathfrak{p}).$
Under the inner product (,)(\ ,\ ), the conormal space $(N^{*}_{L^{\lambda}}W^{\lambda})(g)=\{\phi\in T^{*}_{g}W^{\lambda}\ |\ \phi(T_{g}L^{\lambda})=0\}$ can be identified with $N_{g}:=R_{g*}(L^{+}\mathfrak{g}\cap\mathrm{Ad}_{t^{\lambda}}L^{+}\mathfrak{g}\cap\mathrm{Ad}_{h}L\mathfrak{k})$.
In view of Lemma 15, to show the coisotropicity, we need to show that the contraction of conormal bundle with bivector $\pi$
$N^{*}_{L^{\lambda}}W^{\lambda}\hookrightarrow T^{*}W^{\lambda}|_{L^{\lambda}}\xrightarrow{\pi}TW^{\lambda}|_{L^{\lambda}}$
lands inside $TL^{\lambda}$. Since $T_{g}L^{\lambda}=\{x\in T_{g}W^{\lambda}\subset T_{g}LG\ |\ (x,N_{g})=0\}$, it remains to show that for any $g=h\theta(h)^{-1}\in L^{\lambda}$ and $x,y\in N_{g}$, $(\pi_{g},a\wedge b)=0$. Write $x=R_{g*}\mathrm{Ad}_{h}a$, $y=R_{g*}\mathrm{Ad}_{h}b$, where $a,b\in L\mathfrak{k}$. Since (,)(\ ,\ ) is $LG$-invariant, we have
$(\pi_{g},a\wedge b)=(R_{g*}r-L_{g*}r,R_{g*}\mathrm{Ad}_{h}(a\wedge b))=(\mathrm{Ad}_{h^{-1}}r-\mathrm{Ad}_{\theta(h^{-1})}r,a\wedge b).$
Recall $\theta(r)=r$. Notice that $\theta$ acts trivially on $a,b\in L\mathfrak{k}$ and (,)(\ ,\ ) is $\theta$-invariant. Thus
$(\mathrm{Ad}_{\theta(h^{-1})}r,a\wedge b)=(\mathrm{Ad}_{h^{-1}}r,a\wedge b),$
which completes the proof. ∎
4.6. Lusztig embedding
Let $G=\mathrm{GL}_{n}$. Consider the classical groups $K=\mathrm{O}_{n}$ and $\mathrm{Sp}_{n}$(when $n$ is even) of orthogonal group and symplectic group corresponding to the involution $\theta(g)=(g^{t})^{-1}$ and $\theta(g)=J^{-1}(g^{t})^{-1}J$ respectively. Here $J=\begin{pmatrix}0&I_{m}\\ -I_{m}&0\end{pmatrix}$ with $n=2m$. We have $\mathfrak{g}=\mathfrak{gl}_{n}$, $d\theta(g)=-g^{t}$ (resp. $d\theta(g)=-J^{-1}g^{t}J$), $\mathfrak{k}=\mathfrak{g}^{d\theta}=\mathfrak{o}_{n}$ (resp. $\mathfrak{k}=\mathfrak{sp}_{n}$), $\mathfrak{p}=\mathfrak{g}^{-d\theta}=\{g\in\mathfrak{gl}_{n}\ |\ g^{t}=g\}$ (resp. $\mathfrak{p}=\{g\in\mathfrak{gl}_{n}\ |\ (gJ)^{t}=-gJ\}$).
Let $\mathcal{N}\subset\mathfrak{g}$ be the nilpotent cone and $\mathcal{N}_{\mathfrak{p}}=\mathcal{N}\cap\mathfrak{p}$ be the corresponding symmetric nilpotent cone. The $G$-orbits on $\mathcal{N}$, called the nilpotent orbits, are parametrized by partitions $\mathcal{P}_{n}=\{(\lambda_{1}\geq\lambda_{2}\geq\cdot\cdot\cdot\geq\lambda_{n}\geq 0|\sum\lambda_{i}=n\}$. Let $K^{\circ}\subset K$ be the identity component of $K$. The $K^{\circ}$-orbits on $\mathcal{N}_{\mathfrak{p}}$, called the symmetric nilpotent orbits, are parametrized by $\mathcal{P}_{n}$ in the case of $K^{\circ}=\mathrm{SO}_{n}$ and $\mathcal{P}_{m}$ in the case of $K^{\circ}=K=\mathrm{Sp}_{2m}$. We will write $\mathcal{O}_{\lambda}$ and $\mathcal{O}_{\mathfrak{p},\lambda}$ for the corresponding nilpotent orbits and symmetric nilpotent orbits and $\overline{\mathcal{O}}_{\lambda}$ and $\overline{\mathcal{O}}_{\mathfrak{p},\lambda}$ the orbit closures.
We will identify the set $\Lambda_{T}^{+}$ with $\Lambda_{T}^{+}=\{(\lambda_{1}\geq\lambda_{2}\geq\cdot\cdot\cdot\geq\lambda_{n})\}$ so that
$LG_{\lambda}=G(\mathcal{O})\operatorname{diag}(t^{\lambda_{1}-1},...,t^{\lambda_{n}-1})G(\mathcal{O})$
for $\lambda\in\Lambda_{T}^{+}$. Recall the Lusztig embedding
(7) $\iota:\mathcal{N}\hookrightarrow W^{0}=L^{<0}G,\quad x\mapsto(I_{n}-t^{-1}x).$
It induces an isomorphism
$\mathcal{N}\cong W^{0}_{\leq\omega_{n}}=L^{<0}G\cap\overline{LG}_{\omega_{n}}$
where $\omega_{n}=(n\geq 0\geq\cdots\geq 0)$. In addition, for $\lambda\in\mathcal{P}_{n}$, the above isomorphism restricts to isomorphisms
$\mathcal{O}_{\lambda}\cong W^{0}_{\lambda},\ \ \ \ \ \ \ \overline{\mathcal{O}}_{\lambda}\cong W^{0}_{\leq\lambda}.$
Note that the Lusztig embedding intertwines the involution $d\theta$ on $\mathcal{N}$ with $\operatorname{inv}\circ\theta$ on $W^{0}$ and hence induces isomorphisms
$\mathcal{N}_{\mathfrak{p}}\cong(W^{0}_{\leq\omega_{n}})^{\operatorname{inv}\circ\theta}=L^{0}_{\leq\omega_{n}},$
$\mathcal{O}_{\mathfrak{p},\lambda}\cong(W^{0}_{\lambda})^{\operatorname{inv}\circ\theta}=L^{0}_{\lambda},\ \ \ \ \ \overline{\mathcal{O}}_{\mathfrak{p},\lambda}\cong(W^{0}_{\leq\lambda})^{\operatorname{inv}\circ\theta}=L^{0}_{\leq\lambda}$
on the fixed points. Now in view of Proposition 9, we obtain
Proposition 17.
  • (i)
    There is a $\mathbb{G}_{m}$-equivariant isomorphism
    $\overline{\mathcal{O}}_{\mathfrak{p},\lambda}\cong S^{0}_{\leq\lambda}$
    between symmetric nilpotent orbit closures and transversal slices in loop spaces of classical symmetric varieties.
  • (ii)
    For any $\lambda\leq\mu\leq\omega_{n}\in\mathcal{P}_{n}$, the variety $S^{\lambda}_{\leq\mu}$ defines a Slodowy slice to the symmetric nilpotent orbit $\mathcal{O}_{\mathfrak{p},\lambda}$ inside $\overline{\mathcal{O}}_{\mathfrak{p},\mu}$.
Remark 18.
Specializing Theorem 16 to case in Proposition 17, we obtain (special case of) a well-known result of Kostant-Rallis saying that the symmetric nilpotent orbits $\mathcal{O}_{\mathfrak{p},\lambda}$ are Lagrangian subvarieties of the nilpotent orbits $\mathcal{O}_{\lambda}$ with respect to the Kostant-Kirillov symplectic forms.
For $X=\mathrm{GL}_{2}/\mathrm{O}_{2}$, the slice $S^{0}_{\leq\omega_{2}}$ is isomorphic to $S^{0}_{\leq\omega_{2}}\cong\mathcal{N}_{\mathfrak{p}}\cong\{xy=0\}\subset k^{2}$, which is reducible and non-normal.
For the symmetric variety $\mathrm{GL}_{n}/(\mathrm{GL}_{m}\times\mathrm{GL}_{n-m})$, the involution is given by $\theta(g)=hgh^{-1}$ where $h=\begin{pmatrix}I_{m}&0\\ 0&I_{n-m}\end{pmatrix}$. Then a direct computation shows that the Lusztig embedding does not interwine the involution $d\theta$ and $\operatorname{inv}\circ\theta$ and hence we can’t not conclude Proposition 9 in this case.
5. Formality
The formality conjecture in relative Langlands duality, in the case of symmetric varieties, says that the dg extension algebras for the Satake category for symmetric varieties are formal. We prove the formality conjecture for classical symmetric varieties. The transversal slices constructed in the previous section is a key ingredient in the proof. As an application, we show that the dg extension algebras for the Satake category for the real and quaternionic linear groups $\operatorname{GL}_{n}(\mathbb{R})$ and $\operatorname{GL}_{n}(\mathbb{H})$ are formal.
5.1. Formality conjecture
Let $D^{G(\mathcal{O})}(\mathrm{Gr}_{G})$ be the Satake category for $G$ and let $D^{G(\mathcal{O})}(LX)$ be the dg derived category of $G(\mathcal{O})$-equivariant complexes on $LX$, to be called the Satake category for the symmetric variety $X$. Here we use the !!-sheaves theory, see, e.g., [BZSV, Section 7.2.3]. We have the Hecke action, denoted by $\star$, of the Satake category $D^{G(\mathcal{O})}(\mathrm{Gr}_{G})$ on $D^{G(\mathcal{O})}(LX)$ defined as follows. Consider the maps
$\mathrm{Gr}\times LX\stackrel{{\scriptstyle p\times\operatorname{id}}}{{\leftarrow}}G(F)\times LX\stackrel{{\scriptstyle q}}{{\to}}G(F)\times^{G(\mathcal{O})}LX\stackrel{{\scriptstyle m}}{{\to}}LX$
where $p$ and $q$ are the natural quotient maps and $m$ is the multiplication map. Given $\mathcal{F}_{1}\in D^{G(\mathcal{O})}(\mathrm{Gr})$ and $\mathcal{F}_{2}\in D^{G(\mathcal{O})}(LX)$ we have
(8) $\mathcal{F}_{1}\star\mathcal{F}_{2}:=m_{!}(\mathcal{F}_{1}\tilde{\boxtimes}\mathcal{F}_{2})$
where $\mathcal{F}_{1}\tilde{\boxtimes}\mathcal{F}_{2}\in D^{G(\mathcal{O})}(G(F)\times^{G(\mathcal{O})}LX)$ is the unique complex such that $({p\times\operatorname{id}})^{!}(\mathcal{F}_{1}\boxtimes\mathcal{F}_{2})\cong q^{!}(\mathcal{F}_{1}\tilde{\boxtimes}\mathcal{F}_{2})$. We have the monoidal abelian Satake equivalence
$\operatorname{Rep}(\check{G})\cong\operatorname{Perv}^{G(\mathcal{O})}(\mathrm{Gr}_{G}):\ \ \ V\to\ \mathrm{IC}_{V}.$
By restricting the action to $\operatorname{Perv}^{G(\mathcal{O})}(\mathrm{Gr}_{G})\cong\operatorname{Rep}(\check{G})$, we obtain a monoidal action of $\operatorname{Rep}(\check{G})$ on $D^{G(\mathcal{O})}(LX)$. Let $\omega_{L^{+}X}\in D^{G(\mathcal{O})}(LX)$ be the dualizing complex on the closed $G(\mathcal{O})$-orbit $LX_{0}=L^{+}X$.
Let $\mathrm{IC}_{\mathcal{O}(\check{G})}$ (an ind-object in $\operatorname{Perv}^{G(\mathcal{O})}(\mathrm{Gr}_{G})$) be the image of the regular representation $\mathcal{O}(\check{G})$ under the abelian Satake equivalence. Since $\mathcal{O}(\check{G})$ is a ring object in $\operatorname{Rep}(\check{G})$, the RHom space
$A_{X}:=R\mathrm{Hom}_{D^{G(\mathcal{O})}(LX)}(\omega_{L^{+}X},\mathrm{IC}_{\mathcal{O}(\check{G})}\star\omega_{L^{+}X})$
is naturally a dg-algebra, known as the the de-equivariantized Ext algebra. The formality conjecture is the following assertion (see, e.g., [BZSV, Conjecture 8.1.8]):
Conjecture 19.
The dg-algebra $A_{X}=R\mathrm{Hom}_{D^{G(\mathcal{O})}(LX)}(\omega_{L^{+}X},\mathrm{IC}_{\mathcal{O}(\check{G})}\star\omega_{L^{+}X})$ is formal, that is, it is quasi-isomorphic to the graded algebra $\operatorname{Ext}^{\bullet}_{D^{G(\mathcal{O})}(LX)}(\omega_{L^{+}X},\mathrm{IC}_{\mathcal{O}(\check{G})}\star\omega_{L^{+}X})$ with trivial differential.
We shall show that the formality conjecture holds true when $X$ are classical symmetric varieties.
Remark 20.
Note that $LX_{\mathcal{L}}\subset LX$ is $LG$-invariant and $L^{+}X\subset LX_{\mathcal{L}}$. It follows that $\mathrm{IC}_{V}\star\omega_{L^{+}X}$ lands in the full-subcategory $D^{G(\mathcal{O})}(LX_{\mathcal{L}})\subset D^{G(\mathcal{O})}(LX)$ and we have an isomorphism $A_{X}\cong R\mathrm{Hom}_{D^{G(\mathcal{O})}(LX_{\mathcal{L}})}(\omega_{L^{+}X},\mathrm{IC}_{\mathcal{O}(\check{G})}\star\omega_{L^{+}X})$.
5.2. Placid ind-schemes and dimension theory
We discuss the definition of placid (ind-)schemes and dimension theory following [BKV, Part 1].
Let $Y$ be a scheme acted on by an affine group scheme $H$. we say that $Y$ is $H$-placid if
  • $Y$ can be written as filtered limit $Y=\lim_{j}Y_{j}$, where each $Y_{j}$ is a $H$-scheme of finite type and the transition maps $Y_{j^{\prime}}\to Y_{j}$ for $j\to j^{\prime}$ are affine, smooth, surjective, and $H$-equivariant.
  • the action $H\times Y_{j}\to Y_{j}$ factors through a group scheme $H_{j}$ of finite type. Moreover the $H_{j}$ can be chosen so that $\{H_{j}\}_{j}$ forms a projective system with $H=\operatorname{lim}_{j}H_{j}$ and the transition maps $H_{j^{\prime}}\to H_{j}$ are smooth surjective with unipotent kernel.
Let $Z\subset Y$ a $H$-invariant subscheme. We shall say that the inclusion $Z\to Y$ is placid if there is a presentation of $Y=\operatorname{lim}Y_{j}$ as above and and index $j$ and a $H$-invariant subscheme $Z_{j}\subset Y_{j}$ such that $Z\cong Z_{j}\times_{Y_{j}}Y$.
Let $Y$ be an ind-scheme acted on by an affine group scheme $H$. We say that $Y$ is $H$-placid if can be written as filtered colimit
(9) $Y\cong\operatorname{colim}_{i}\lim Y^{i}$
where each $Y^{i}$ is placid and the transition maps $Y^{i}\to Y^{i^{\prime}}$ are placid closed embeddings. A presentation as in (9) is called a placid presentation of $Y$. If $Y$ is $H$-placid, then the category of placid presentations of $Y$ has an initial object, and in particular is contractible. So any two placid presentations are essentially equivalent.
Let $Y$ be a placid ind-scheme as in (9). We have a function $\dim_{Y^{i}_{j}/Y^{i^{\prime}}_{j}}:Y^{i}_{j}\to\mathbb{Z}$ given by $\dim_{Y^{i}_{j}/Y^{i^{\prime}}_{j}}(y)=\dim_{y}(Y^{i}_{j})-\dim_{y}(Y^{i^{\prime}}_{j})$. Here for a finite type scheme $U$ we denote by $\dim_{x}(U)$ the maximal of dimensions of irreducible components containing $x$. According to [BKV, Lemma 2.2.4-2.2.5] there exists a unique function $\dim_{Y^{i}/Y^{i^{\prime}}}:Y^{i}\to\mathbb{Z}$ such that $\dim_{Y^{i}/Y^{i^{\prime}}}=(p^{i}_{j})^{*}(\dim_{Y^{i}_{j}/Y^{i^{\prime}}_{j}})$ where $p^{i}_{j}:Y^{i}\to Y^{i}_{j}$ is the projection map. A dimension theory $\delta$ on $Y$ is an assignment to each $i$ a locally constant function $\delta_{Y^{i}}:Y^{i}\to\mathbb{Z}$ such that $\delta_{Y^{i}}=\delta_{Y^{i^{\prime}}}|_{Y^{i}}+\dim_{Y^{i}/Y^{i^{\prime}}}$.
5.3. Sheaf theories
We will be working with $\mathbb{C}$-linear dg-categories. Unless specified otherwise, all dg-categories will be assumed cocomplete, i.e., containing all small colimits, and all functors between dg-categories will be assumed continuous, i.e., preserving all small colimits. For any finite type scheme $Y$ acted on by an affine algebraic group $H$ we denote by $D^{H}(Y)$ the ind-completion of the the bounded equivariant dg derived category.
For any $H$-placid ind-scheme $Y$, we define $D^{H}_{?}(Y)=\operatorname{colim}_{i,!}\operatorname{colim}_{j}^{?}D^{H^{i}_{j}}(Y^{i}_{j})$ where the first colimit is with respect to the !!-pushforward along the closed embedding $Y^{i}\to Y^{i^{\prime}}$ and the second colimit is with respect to ??-pull back along the map $Y^{i}_{j}/H^{i}_{j}\to Y^{i}_{j^{\prime}}/H^{i}_{j^{\prime}}$ where ${!,,!}?\in\{$!,*,!*\}. Since !!*-pull back is $t$-exact with respect to the perverse $t$-structures on $D^{H^{i}_{j}}(Y^{i}_{j})$, the category $D_{!*}^{H}(Y)$ carries a canonical perverse $t$-structure and we denote by $\operatorname{Perv}^{H}(Y)$ the corresponding category of perverse sheaves.
Let $Z\subset Y$ be a $H$-invariant closed subscheme. Assume the embedding $f:Z\to Y$ is placid then we have the $\mathrm{IC}$-complex on $Z$
$\mathrm{IC}_{Z}=\operatorname{colim}_{j}(f_{j})_{*}\mathrm{IC}_{Z_{j}}\in\operatorname{Perv}^{H}(Y)$
where $\mathrm{IC}_{Z_{j}}$ are the $\mathrm{IC}$-complexes on $Z_{j}$ and $f_{j}:Z_{j}\to Y_{j}$ is the closed embedding.
We will need the following lemma in Section 5.4.
Lemma 21.
Let $Y$ be a $H$-placid ind-scheme. If there exits a dimension theory $\delta$ on $Y$, then we have equivalences
$D^{H}_{!}(Y)\cong D^{H}_{!*}(Y)\cong D^{H}_{*}(Y).$
Proof.
From the definition of the categories $D^{G(\mathcal{O})}_{?}(Y)$, the assignment sending $\{\mathcal{F}^{i}_{j}\}\in D_{!}^{H}(Y)$, to $\{\mathcal{F}^{i}_{j}[\dim(Y^{i}_{j})-\delta_{Y^{i}}]\}\in D_{!*}^{H}(Y)$ (resp. $\{\mathcal{F}^{i}_{j}[2\dim(Y^{i}_{j})-2\delta_{Y^{i}}]\}\in D_{*}^{H}(Y)$ ) defines the desired equivalence $D^{G(\mathcal{O})}_{!}(Y)\cong D^{G(\mathcal{O})}_{!*}(Y)$ (resp. $D^{G(\mathcal{O})}_{!*}(Y)\cong D^{G(\mathcal{O})}_{*}(Y)$). ∎
5.4. Classical symmetric varieties
Let $X$ be the classical symmetric variety $X=\operatorname{GL}_{n}/\operatorname{O}_{n}$ or $\operatorname{GL}_{2m}/\operatorname{Sp}_{2m}$ as in Section 4.6. Since $\operatorname{Sp}_{2m}$ is connected, we have $\mathcal{L}=\Lambda_{S}^{+}$ and $LX=LX_{\mathcal{L}}$ in the case $X=\operatorname{GL}_{2m}/\operatorname{Sp}_{2m}$.
Proposition 22.
  • (i)
    LXLX is a $G(\mathcal{O})$-placid ind-scheme.
  • (ii)
    $\overline{LX}_{\lambda}$ are $G(\mathcal{O})$-placid schemes and the inclusion $\overline{LX}_{\lambda}\to LX$ is placid.
  • (iii)
    $LX_{\mathcal{L}}$ is $G(\mathcal{O})$-placid and there exist a dimension theory on $LX_{\mathcal{L}}$.
Proof.
Recall the isomorphism $\tau:X\cong G^{\operatorname{inv}\circ\theta,\circ}$. When $X$ is a classical symmetric variety, we have $G^{\operatorname{inv}\circ\theta,\circ}=\mathring{\mathfrak{p}}\subset\mathfrak{p}$, where $\mathring{\mathfrak{p}}$ is the open subset consisting of invertible matrices in $\mathfrak{p}$. Hence for any $k$-algebra $R$, we can and will identify the $R$-points $LX(R)$ with points $\gamma\in L\mathfrak{p}(R)=\mathfrak{p}(R((t)))$ satisfying $\operatorname{det}(\gamma)\in R((t))^{\times}$.
Proof of (i). For any $l\in\mathbb{Z}$, let $LX_{l}=\{\gamma\in LX|\operatorname{det}(\gamma)=ct^{l}\text{\ for some\ }c\in\mathcal{O}^{\times}\}$. Note that each $LX_{l}$ is a union of connected components of $LX$ and is $G(\mathcal{O})$-invariant. For any $i\in\mathbb{Z}$, we define $L^{i}X_{l}$ to be following fiber product
${L^{i}X_{l}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${LX_{l}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${t^{i}L^{+}\mathfrak{p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${L\mathfrak{p}}$
The transition maps $t^{i}L^{\mathfrak{p}}\to t^{i^{\prime}}L\mathfrak{p}$ induced closed embeddings $L^{i}X_{l}\to L^{i^{\prime}}X_{l}$ and we have $LX_{l}=\operatorname{colim}_{i}L^{i}X_{l}$. Note that for any $\gamma\in L^{i}X_{l}$ there exists $j\gg i$ such that for any $M\in L^{+}\mathfrak{p}$ we have $\operatorname{det}(\gamma+t^{j}M)=ct^{l}$ for some $c\in\mathcal{O}^{\times}$. Thus we have a translation action of $t^{j}L^{+}\mathfrak{p}$ on $L^{i}X_{l}$ for $j\gg i$ and the quotient $L^{i}_{j}X_{l}$ is of finite type and the action of $G(\mathcal{O})$ on $L^{i}_{j}X_{l}$ factors through a congruence subgroup. Thus we obtain a presentation $LX_{l}=\operatorname{colim}_{i}\operatorname{lim}_{j}L^{i}_{j}X_{l}$ which is $G(\mathcal{O})$-placid
Proof of (ii). The proof is similar. We first show that for any orbit closure $\overline{LX}_{\lambda}$, there exists an integer $j\gg 0$ such that $\overline{LX}_{\lambda}+t^{j}L^{+}\mathfrak{p}=\overline{LX}_{\lambda}$, Note that $LX_{\lambda}(R)=\mathfrak{p}(R((t)))\cap G(R[[t]])t^{\lambda}G(R[[t]])$. Since the lattice $t^{h}L^{+}\mathfrak{p}(R)=t^{h}\mathfrak{p}(R[[t]])$ is $G(R[[t]])$-invariant, it suffices to show $t^{\lambda}+t^{h}\mathfrak{p}(R[[t]])\in\mathfrak{p}(R((t)))$ lies in the coset $G(R[[t]])t^{\lambda}G(R[[t]])$. This follows from the fact that
$t^{\lambda}+t^{h}\mathfrak{p}(R[[t]])=t^{\lambda}(\operatorname{Id}+t^{-\lambda}t^{h}\mathfrak{p}(R[[t]]))$
and
$(\operatorname{Id}+t^{-\lambda}t^{h}\mathfrak{p}(R[[t]]))\subset G(R[[t]])$
for $h$ large enough with respect to $\lambda$. Since $\overline{LX}_{\lambda}=\bigsqcup_{\lambda^{\prime}\leq\lambda}LX_{\lambda^{\prime}}$ is a finite union of orbits, we can take $j$ be the maximal one among that for all the $LX_{\lambda^{\prime}}$.
Since $\overline{LX}_{\lambda}\subset L^{i}X_{l}$ for some $l$ and large $i$ we see that the closed embedding $\overline{LX}_{\lambda}\to L^{i}X_{l}$ is the limit of the closed embedding $\overline{LX}_{\lambda}/t^{j}L^{+}\mathfrak{p}\to L^{i}X_{l}/t^{j}L^{+}\mathfrak{p}=L^{i}_{j}X_{l}$ and it implies that $\overline{LX}_{\lambda}$ is $G(\mathcal{O})$-placid and the closed embedding $\overline{LX}_{\lambda}\to LX$ is placid.
Proof of (iii). It suffices to show $LX_{\mathcal{L},l}=LX_{\mathcal{L}}\cap LX_{l}$ is $G(\mathcal{O})$-placid and has a dimension theory. Let $L^{i}X_{\mathcal{L},l}:=L^{i}X_{l}\cap LX_{\mathcal{L}}$. Note that there are only finitely many $G(\mathcal{O})$-orbits on $L^{i}X_{l}\cap LX_{\mathcal{L}}$ and part (ii) implies that $L^{i}X_{\mathcal{L},l}+t^{j}L^{+}\mathfrak{p}=L^{i}X_{\mathcal{L},l}$ for $j\gg i$. Let $L^{i}_{j}X_{\mathcal{L},l}=L^{i}X_{\mathcal{L},l}/t^{j}L^{+}\mathfrak{p}$ and we get a placid presentation $LX_{\mathcal{L},l}=\operatorname{colim}_{i}\operatorname{lim}_{j}L_{j}^{i}X_{\mathcal{L},l}$ For any $x\in L^{i}X_{\mathcal{L},l}$ let $LX_{\lambda}$ be an open orbit in $L^{i}X_{\mathcal{L},l}$ such that $x\in\overline{LX}_{\lambda}$ and $\langle\lambda,\rho\rangle\in\frac{1}{2}\mathbb{Z}$ is maximal along all open orbits, and we define $\delta_{L^{i}X_{\mathcal{L},l}}(x)=\langle\lambda,\rho\rangle$ if $\langle\lambda,\rho\rangle\in\mathbb{Z}$ or $\delta_{L^{i}X_{\mathcal{L},l}}(x)=\langle\lambda,\rho\rangle+1/2$ if $\langle\lambda,\rho\rangle\in\frac{1}{2}\mathbb{Z}$. Then it follow from Corollary 14 that $\delta_{L^{i}X_{\mathcal{L},l}}:L^{i}X_{\mathcal{L},l}\to\mathbb{Z},x\to\delta_{L^{i}X_{\mathcal{L},l}}(x)$ defines a dimension theory for $LX_{\mathcal{L},l}$. ∎
Theorem 23.
The extension algebra $A_{X}$ is formal.
Proof.
Recall that $A_{X}\cong R\mathrm{Hom}_{D^{G(\mathcal{O})}_{!*}(LX_{\mathcal{L}})}(\omega_{L^{+}X},\mathrm{IC}_{\mathcal{O}(\check{G})}\star\omega_{L^{+}X})$, see Remark 20. By Lemma 21 and Proposition 22.(iii), there exists an equivalence $D^{G(\mathcal{O})}_{!}(LX_{\mathcal{L}})\cong D^{G(\mathcal{O})}_{!*}(LX_{\mathcal{L}})$. We also have the Hecke action of $D^{G(\mathcal{O})}(\mathrm{Gr})$ on $D^{G(\mathcal{O})}_{!*}(LX_{\mathcal{L}})$ defined by the same formula as in (8) and it follows from the definition that the equivalence above is compatible with the Hecke action. Write $\mathrm{IC}_{\lambda}=\mathrm{IC}_{\overline{LX}_{\lambda}}\in\operatorname{Perv}^{G(\mathcal{O})}(LX_{\mathcal{L}})$ for the $\mathrm{IC}$-complexes for the orbit closures $\overline{LX}_{\lambda}$ and
$B_{X}:=R\mathrm{Hom}_{D^{G(\mathcal{O})}_{!*}(LX_{\mathcal{L}})}(\mathrm{IC}_{0},\mathrm{IC}_{\mathcal{O}(\check{G})}\star\mathrm{IC}_{0})$
for the de-equivariantized Ext algebra for $D^{G(\mathcal{O})}_{!*}(LX_{\mathcal{L}})$. Then we have an isomorphism of dg-algebra $A_{X}\cong B_{X}$ and we reduce to show that $B_{X}$ is formal. It follows from the equivalence in (10) and [NadlerRealGr, Lemma 10.5.1] that $\mathrm{IC}_{\mathcal{O}(\check{G})}\star\mathrm{IC}_{0}$ is a direct sum of shifted $\mathrm{IC}$-complexes $\mathrm{IC}_{\lambda}$ (with constant coefficients). Then by the result of [BF, Section 6.5], the formality of $B_{X}$ follows from the the purity of $\operatorname{Ext}^{\bullet}_{D^{G(\mathcal{O})}_{!*}(LX_{\mathcal{L}})}(\mathrm{IC}_{0},\mathrm{IC}_{\lambda})$. Since $\operatorname{Ext}^{\bullet}_{D^{G(\mathcal{O})}_{!*}(LX_{\mathcal{L}})}(\mathrm{IC}_{0},\mathrm{IC}_{\lambda})\cong H^{\bullet}_{G}(\operatorname{pt})\otimes i^{!}_{0}\mathrm{IC}_{\lambda}$, where $i_{0}:\{x_{0}\}\to LX_{\mathcal{L}}$ is the base point, it suffices to show the purity of the stalk $i^{!}_{0}\mathrm{IC}_{\lambda}$. Since $\mathrm{IC}_{\lambda}$ is the !!*-pullback of $\mathrm{IC}$-complex on the quotient $Z:=\overline{LX}_{\lambda}/t^{j}L^{+}\mathfrak{p}$ (which is of finite type) for some $j\gg 0$, it suffices to show the purity of the stalk of $\mathrm{IC}_{Z}$ at the base point $z_{0}$, the image of $x_{0}$ in $Z$. Consider the transversal slice $S^{0}_{\leq\lambda}\subset\overline{LX}_{\lambda}$ at $x_{0}$. Note that the composition $S^{0}_{\leq\lambda}\subset\overline{LX}_{\lambda}\to Z=\overline{LX}_{\lambda}/t^{j}L^{+}\mathfrak{p}$ is an injective for $j\gg 0$ and defines a transversal slice at $z_{0}$ of $Z$. Since the $\mathrm{IC}$-complex $\mathrm{IC}_{Z}$ is pure of weight zero and $\mathbb{G}_{m}$-equivariant we can conclude the proof by applying the general fact in [MS, Proposition 2.3.3]. ∎
5.5. Real groups
Let $G_{\mathbb{R}}$ be the real from of $G$ corresponding to $K$. Let $F_{\mathbb{R}}=\mathbb{R}((t))$ and $\mathcal{O}_{\mathbb{R}}=\mathbb{R}[[t]]$. Let $G_{\mathbb{R}}(F_{\mathbb{R}})$ and $G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})$ be the real loop and arc groups and let $\mathrm{Gr}_{\mathbb{R}}=G_{\mathbb{R}}(F_{\mathbb{R}})/G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})$ be the real affine Grassmannian. Consider the dg category $D^{G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})}(\mathrm{Gr}_{\mathbb{R}})$ of $G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})$-equivariant complexes on the real affine Grassmannian $\mathrm{Gr}_{\mathbb{R}}$ (with respect to the classical topology). It is shown in [NadlerRealGr] that there is a monoidal functor $\psi:D^{G(\mathcal{O})}(\mathrm{Gr})\to D^{G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})}(\mathrm{Gr}_{\mathbb{R}})$ given by the nearby cycle functor assoicated to a real form of Beilinson-Drinfeld Grassmannian. Thus we have a natural monoidal action of $\operatorname{Rep}(G)\cong\operatorname{Perv}^{G(\mathcal{O})}(\mathrm{Gr})$ on $D^{G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})}(\mathrm{Gr}_{\mathbb{R}})$, denoted by $\star$, and we can form the de-equivariantized Ext algebra
$A_{\mathbb{R}}:=R\mathrm{Hom}_{D^{G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})}(\mathrm{Gr}_{\mathbb{R}})}(\delta_{\mathbb{R}},\mathrm{IC}_{\mathcal{O}(\check{G})}\star\delta_{\mathbb{R}})$
where $\delta_{\mathbb{R}}$ is the delta sheaf supported at the base point of $\mathrm{Gr}_{\mathbb{R}}$.
We have the following result.
Proposition 24.
There is an isomorphism of dg-algebra $A_{X}\cong A_{\mathbb{R}}$.
Proof.
It is shown in [NadlerRealGr, Lemma 3.9.1] that the union of real spherical orbits $\mathrm{Gr}_{\mathbb{R},\mathcal{L}}:=\bigcup_{\lambda\in\mathcal{L}}\mathrm{Gr}_{\mathbb{R}}^{\lambda}$ is a union of connected components of $\mathrm{Gr}_{\mathbb{R}}$. Moreover, there is a real and symmetric equivalence [CN1, Theorem 8.1]
(10) $D^{G(\mathcal{O})}(LX_{\mathcal{L}})\cong D^{G(\mathcal{O})}(LG/LK)\cong D^{G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})}(\mathrm{Gr}_{\mathbb{R},\mathcal{L}})$
which is compatible with the action of the Satake category $D^{G(\mathcal{O})}(\mathrm{Gr})$ and maps $\omega_{L^{+}X}\in D^{G(\mathcal{O})}(LX_{\mathcal{L}})$ to $\delta_{\mathbb{R}}\in D^{G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})}(\mathrm{Gr}_{\mathbb{R},\mathcal{L}})$. Similar to Remark 20, we have
$A_{\mathbb{R}}\cong R\mathrm{Hom}_{D^{G_{\mathbb{R}}(\mathcal{O}_{\mathbb{R}})}(\mathrm{Gr}_{\mathbb{R},\mathcal{L}})}(\delta_{\mathbb{R}},\mathrm{IC}_{\mathcal{O}(\check{G})}\star\delta_{\mathbb{R}})$
and thus
[…]
Theorem 23 immediately implies:
Corollary 25.
The dg-algebra $A_{\mathbb{R}}$ is formal in the case $G_{\mathbb{R}}=\mathrm{GL}_{n}(\mathbb{R})$ and $\mathrm{GL}_{n}(\mathbb{H})$.
Remark 26.
The formality of $A_{\mathbb{R}}$ for $\mathrm{GL}_{n}(\mathbb{H})$ was proved in [QuaternionicSatake] by a different method. The proof in loc. cit. uses the fact that that the real nearby cycle fucntor induces a surjective map from the dg-extension algebra for the complex Grassmannian $\mathrm{Gr}$ to $A_{\mathbb{R}}$. The approach to formality given here relies on the fact that the loop space $LX$ of classical symmetric varieties $X$ is placid and admits a dimension theory. It would be nice if this fact is true for all symmetric varieties and then one can apply the same argument to get a uniform proof of the formality conjecture in general.